The mean weight of trucks traveling on a particular section I-475 is not known. A state highway inspector needs an estimate of the mean. He selects a random sample of 49 truck passing the weighing station and finds the mean is 15.8 tons. He sample standard deviation is 5.0 tons. What is the 95 percent interval for the population mean?
Answer by stanbon(75887)

The mean weight of trucks traveling on a particular section I-475 is not known. A state highway inspector needs an estimate of the mean. He selects a random sample of 49 truck passing the weighing station and finds the mean is 15.8 tons. He sample standard deviation is 5.0 tons. What is the 95 percent interval for the population mean?
---
sample mean = 15.8
standard error: (invT(0.975 with df = 48)*5/sqrt(49) = 2.011*(5/7) = 1.4362
-----
95% CI: 15.8-1.4362 < u < 15.8+1.4362
14.36 < u < 17.24
=======================
Cheers,
Stan H.
=======================